Suppose you graphed every single point of the form (2t + 3, 3-3t). For example, when t=2, we have 2t + 3 = 7 and 3-3t = -3, so (7,-3) is on the graph. Explain why the graph is a line, and find an equation whose graph is this line.

Answer by stanbon(75887)

Suppose you graphed every single point of the form (2t + 3, 3-3t). For example, when t=2, we have 2t + 3 = 7 and 3-3t = -3, so (7,-3) is on the graph. Explain why the graph is a line, and find an equation whose graph is this line.
-----
(2t+3,3-3t) is the parametric form of a line.
If t = 0, you get the point (3,3)
---
Find the line thru (3,3) and (7,-3)
slope = (3--3)/(3-7) = 6/-4 = -3/2
Form: y = mx + b
3 = (-3/2)3 + b
b = 3+(9/2) = 7.5
----
Equation::
y = (-3/2)x + (15/2)
--------------
Cheers,
Stan H.
---------------------